AN IMPROVED ANALYSIS OF THE SEVOFLURANE-BENZENE STRUCTURE BY CHIRPED PULSE FTMW SPECTROSCOPY

\underline{NATHAN A. SEIFERT}, CRISTOBAL PEREZ, DANIEL P. ZALESKI, JUSTIN L. NEILL, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904-4319; ALBERTO LESARRI, MONTSERRAT VALLEJO, Departamento de Química Física y Química Inorgánica, Facultad de Ciencias, Universidad de Valladolid, E-47011 Valladolid, Spain; EMILIO J. COCINERO, FERNANDO CASTANO, Departamento de Química Física, Facultad de Ciencia y Tecnología, Universidad del País Vasco (UPV/EHU), Campus de Leioa, Ap. 644, E-48080 Bilbao, Spain; ISABELLE KLEINER, Laboratorie Interuniversitaire des Systèmes Atmosphériques (LISA), CNRS et Universités Paris Est et Paris Diderot, 61. av Général de Gaulle, 94010 Créteil, France.

Recent improvements to the 2-8 GHz CP-FTMW spectrometer at University of Virginia have improved the structural and spectroscopic analysis of the sevoflurane-benzene cluster. Previously reported results^{a}, although robust, were limited to a fit of the a-type transitions of the normal species in the determination of the six-fold barrier to benzene internal rotation. Structural analysis was limited to the benzene hydrogen atom positions using benzene-d_{1}. The increased sensitivity of the new 2-8 GHz setup allows for a full internal rotation analysis of the a- and c-type transitions of the normal species, which was performed with BELGI. A fit value for V_{6} of 32.868(11) cm^{−1} is determined. Additionally, a full substitution structure of the benzene carbon atom positions was determined in natural abundance. Also, new measurements of a sevoflurane/benzene-d_{1} mixture enabled detection of 33 of the 60 possible ^{2}D / ^{13}C double isotopologues. This abundance of isotopic data, a total of 45 isotopologues, enabled a full heavy atom least-squares r_{0} structure fit for the complex, including positions for all seven fluorines in sevoflurane.

aN. A. Seifert, D. P. Zaleski, J. L. Neill, B. H. Pate, A. Lesarri, M. Vallejo, E. J. Cocinero, F. Castańo. 67th OSU Int. Symp.

On Mol. Spectrosc., Columbus, OH, 2012, MH13.